OpenStudy (lovelycharm):

Let f(x) = 2x - 6 Solve f-1(x) when x = 2.

OpenStudy (anonymous):

are you trying to solve for f?

OpenStudy (lovelycharm):

yh

OpenStudy (anonymous):

start with f(2)=2(2)

OpenStudy (kc_kennylau):

do you mean f^-1(x) as in \(f^{-1}(x)\)

OpenStudy (lovelycharm):

yh

OpenStudy (anonymous):

Presuming you mean the inverse of f(x), then we can find the inverse by looking at the original function and "undoing" it. We see that the original function adds 6 to 2x. So we take away 6 and divide by 2. f(−1)(x) = (x-6)/2. We can check this easily by putting x=(x-6)/2 into the original f(x). So simply substitute the value you want into the inverse equation.

OpenStudy (lovelycharm):

i still dont get it